# Decoherence-free interaction and maximally entangled state generation in double-giant-atom semi-infinite waveguide systems

### v1

@@ Abstract
Giant atoms are artificial atoms that can couple to a waveguide non-locally. Previous work has shown that two giant atoms in a braided configuration can interact through a one-dimensional infinite waveguide, with both individual and collective atomic relaxation being fully suppressed. In this paper, however, we show that the decoherence-free interaction (DFI) between two giant atoms can be realized in both braided and nested configurations when the waveguide is semi-infinite. This protected interaction fails to appear in semi-infinite waveguide systems containing two separate giant atoms or two small atoms. We also study the entanglement generation between two giant atoms coupled to a one-dimensional semi-infinite waveguide. The results show that the maximally entangled state is generated in both braided and nested configurations due to the formation of DFI, and in the separate configuration, the maximally achievable entanglement can exceed 0.5. This study presents a new scheme for realizing DFI and generating maximally entangled states in giant-atom waveguide-QED systems.
@@ I INTRODUCTION
Recently, waveguide-QED systems containing giant atoms have garnered significant attention from many peers. In such systems, giant atoms are artificial atoms that can interact with the waveguide at multiple connection points [17]. Several interesting quantum optical phenomena have been demonstrated in giant-atom waveguide-QED systems, including frequency-dependent relaxation rates and Lamb shifts [18], non-Markovian decay dynamics [19, 20, 21], single-photon scattering [22, 23, 24], and DFI [25, 26, 27, 28, 29]. In addition, the generation [30, 31, 32, 33, 34] and transfer [35] of quantum entanglement have also been investigated in giant-atom waveguide-QED systems.
However, it is important to note that in most studies on giant atoms, the waveguide is usually considered infinite. A key question is how giant atoms behave when coupled to a semi-infinite waveguide. The semi-infinite waveguide is a one-dimensional waveguide with one end terminated by a perfect mirror [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. Compared to the infinite waveguide, the semi-infinite waveguide is particularly notable in that the mirror at the end allows the atom to interact twice with the field in the waveguide. Hence, when the giant atom couples to the semi-infinite waveguide, its behavior may differ from that in the infinite waveguide. While the semi-infinite waveguide setup containing giant atoms has been addressed in previous studies [18, 49, 50], only a single giant atom was considered.
In this work, our study goes beyond the framework of a single giant atom coupled to a semi-infinite waveguide. We consider an arbitrary number of giant atoms $M$ interacting with a semi-infinite waveguide. Treating the semi-infinite waveguide as the environment of atoms, we derive the Markov quantum master equation governing the dynamical evolution of $M$ giant atoms.
We focus on the case of two giant atoms. Concretely, we analyze the behavior of two giant atoms coupled to a one-dimensional semi-infinite waveguide with three configurations: braided, separate, and nested. In doing so, we show that the DFI between two giant atoms, previously only observed within the braided configuration in infinite waveguide systems [25], can be realized in both braided and nested configurations in semi-infinite waveguide systems. This protected interaction is not supported in a semi-infinite waveguide setup involving either two separate giant atoms or two small atoms. For two small atoms coupled to a semi-infinite waveguide, it has been reported that there is no DFI between them [25].
@@
Figure 1: (Color online) Two giant atoms in a braided configuration coupled to a one-dimensional semi-infinite waveguide. The connection points are located at coordinates $x_{j1}$ and $x_{j2}$, where $j=a,b$. In our work, we consider the case where the two giant atoms share the same transition frequency, i.e., $\omega_{a}=\omega_{b}=\omega_{0}$.
@@
This paper begins with Sec. II, which describes the model for giant atoms coupled to a one-dimensional semi-infinite waveguide. Then, we specifically explore the DFI and entanglement generation between two braided giant atoms in Sec. III and IV. In Sec. V, we consider the case of two giant atoms in the separate and nested configurations. Finally, we discuss the experimental feasibility of the proposed scheme and summarize the main findings of this study in Sec. VI. In addition, we provide two Appendixes: Appendix A and Appendix B. In Appendix A, the detailed derivation of the master equation is shown. In Appendix B, we present the analytical expressions for the concurrence that depend on the phase shift.
@@ II THEORETICAL FRAMEWORK
We start by formulating a general model in which $M$ two-level giant atoms are coupled to a semi-infinite waveguide along the positive $x$ semi-axis, with each atom $j$ having $N_{j}$ connection points. In this model, we assume the semi-infinite waveguide is realized by terminating the left end (coordinate set to $x=0$) of the infinite waveguide with a mirror, and all atoms are located on the right side of the mirror.
@@
From Eq. (2), the primary difference with respect to the infinite waveguide case [25] is that the expressions for $\delta\omega_{j}$, $g_{j,k}$, $\Gamma_{j}$, and $\Gamma_{coll,j,k}$ no longer depend just on the relative distances between the connection points: they are also determined by the distance from the connection points to the mirror in the semi-infinite waveguide case.
We focus on the case of two giant atoms where each atom interacts with the waveguide at two connection points. For two giant atoms with two connection points each, three distinct topology configurations are possible: braided, separate, and nested [25]. The setup we initially consider is the braided configuration, as shown in Fig. 1. Later, we consider the separate and nested configurations in Sec. V. To simplify the discussion, the model in Fig. 1 adopts the following assumptions: $(i)$ the decay rate at each connection point is identical and denoted by $\gamma$, and $(ii)$ the phase shift between adjacent connection points is defined as $\theta$, where $\theta/2$ represents the phase shift accumulated as a photon travels from position $x_{a1}$ to the mirror ($x=0$). When the system contains two giant atoms, Eq. (1) reduces to
@@
$+\Gamma_{coll,a,b}\left[(\sigma_{-}^{a}\rho\sigma_{+}^{b}-\frac{1}{2}\left\{\sigma_{+}^{b}\sigma_{-}^{a},\rho\right\})+\text{H.c.}\right],$
@@
Figure 2: (Color online) Exchange interaction and decay rates as a function of $\theta/\pi$.
III DECOHERENCE-FREE INTERACTION
In previous work, it has been confirmed that two braided giant atoms can interact via the infinite waveguide without losing their excitations to it [25]. That is, there is a decoherence-free interaction (DFI) between two braided giant atoms. Mathematically, the condition for DFI to manifest can be stated as
@@ IV.1 Concurrence
In our work, we focus on entanglement generation in the case where only one giant atom is excited. When there is only one excitation in the system, we can remove the jump term in the quantum master equation (3) to obtain a non-Hermitian effective Hamiltonian $H_{\text{eff}}$ [30, 34, 69, 70]:
$\hat{H}_{\text{eff}}$ $=\sum_{j=a,b}\delta\omega_{j}\hat{\sigma}_{+}^{j}\hat{\sigma}_{-}^{j}+g_{{a,b}}(\hat{\sigma}_{+}^{b}\hat{\sigma}_{-}^{a}+\hat{\sigma}_{+}^{a}\hat{\sigma}_{-}^{b})$ (7)
$\quad-\frac{i}{2}\sum_{j=a,b}\Gamma_{j}\hat{\sigma}_{+}^{j}\hat{\sigma}_{-}^{j}-\frac{i}{2}\Gamma_{coll,a,b}(\hat{\sigma}_{+}^{b}\hat{\sigma}_{-}^{a}+\hat{\sigma}_{+}^{a}\hat{\sigma}_{-}^{b}).$
@@
$i\frac{\partial}{\partial t}|\psi(t)\rangle=\hat{H}_{\text{eff}}|\psi(t)\rangle,$ (8)
@@ IV.2 Entanglement generation
$C_{B}(t)=\frac{e^{-\frac{1}{2}\,\mathrm{Re}\left[R_{1}(t)\right]}}{2}\left|\frac{R_{2}(t)\,\left(R_{3}(t)+R_{4}(t)\right)}{AA^{*}}\right|.$ (12)
where the subscript $B$ denotes the braided configuration case. Similarly, subscripts $S$ and $N$, used in Sec. V, correspond to the separate and nested configurations, respectively. The expressions for $A$ and $R_{i}(t)$ ($i=1,2,3,4$), which depend on the phase shift $\theta$, are given in Appendix B.
@@
To clearly elucidate the phase-shift dependence of entanglement generation, we now analyze the evolution of the concurrence $C_{B}$ when phase shift $\theta$ takes some specific values, where $\theta\in[0,\pi]$. As shown in Fig. 3(a)(a), when $\theta=0$, the concurrence $C_{B}$ exhibits steady-state behaviors.
Figure 3: (Color online) $(a)$ Concurrence $C_{B}$ as a function of $\theta/\pi$ and $\gamma t$. $(b)$ Concurrence $C_{B}(t)$ as a function of $\gamma t$ at a given value of phase shift.
The steady-state behavior of the concurrence $C_{B}$ at $\theta=0$ can be explained through the analytical expression. When $\theta=0$, the concurrence in Eq. (12) becomes
$C_{B}(t)=\frac{1}{2}\left({1-e^{-16\gamma t}}\right).$ (13)
@@
$C_{B}(t)=|\sin(2\gamma t)|.$ (14)
According to Eq. (14), it can be found that the concurrence $C_{B}(t)$ oscillates between zero and one with a period $\frac{\pi}{2\gamma}$. When $t=\frac{n\pi}{4\gamma}$ (where $n$ is an odd integer), concurrence $C_{B}(t)=1$, indicating that the two giant atoms reach a maximally entangled state. Specifically, when $\theta$ is taken to be $\pi$, no entanglement is generated here, i.e., $C_{B}=0$. The reason is mainly that when $\theta=\pi$, exchange interaction $g_{a,b}$, individual decay rates $\Gamma_{a}$ and $\Gamma_{b}$, as well as the collective decay rate $\Gamma_{\text{coll},a,b}$ all become zero (as can be seen from Fig. 2), which implies that the two giant atoms decouple from the waveguide, thereby making entanglement generation impossible.
Figure 4: (Color online) $(a)$ Two giant atoms in a separate configuration coupled to a one-dimensional semi-infinite waveguide. $(b)$ Exchange interaction and decay rates as a function of $\theta/\pi$. $(c)$ Concurrence $C_{s}$ as a function of $\theta/\pi$ and $\gamma t$.
Figure 5: (Color online) $(a)$ Two giant atoms in a nested configuration coupled to a one-dimensional semi-infinite waveguide. $(b)$ Exchange interaction and decay rates as a function of $\theta/\pi$. $(c)$ Concurrence $C_{N}$ as a function of $\theta/\pi$ and $\gamma t$.
@@ V TWO GIANT ATOMS IN SEPARATE AND NESTED CONFIGURATIONS
In the above sections, we have analyzed the behavior of two braided giant atoms coupled to a semi-infinite waveguide, as well as their entanglement properties. In this section, as shown in Fig. 4 and Fig. 5, we consider the separate and nested configurations. According to Eq. (2), the relevant parameters for separate [Fig. 4(a)(a)] and nested [Fig. 5(a)(a)] configurations can be obtained as
$\delta\omega_{a}$ $=\gamma\sin\theta+\gamma\sin(2\theta)+\frac{\gamma}{2}\sin\theta+\frac{\gamma}{2}\sin(3\theta),$ (15)
$\delta\omega_{b}$ $=\gamma\sin\theta+\gamma\sin(6\theta)+\frac{\gamma}{2}\sin(5\theta)+\frac{\gamma}{2}\sin(7\theta),$
$g_{a,b}$ $=\frac{\gamma}{2}(\sin\theta+2\sin(2\theta)+2\sin(3\theta)+2\sin(4\theta)$
$\quad+\sin(5\theta)),$
$\Gamma_{a}$ $=2\gamma+2\gamma\cos\theta+\gamma\cos\theta+\gamma\cos(3\theta)$
$\quad+2\gamma\cos(2\theta),$
$\Gamma_{b}$ $=2\gamma+2\gamma\cos\theta+\gamma\cos(5\theta)+\gamma\cos(7\theta)$
$\quad+2\gamma\cos(6\theta),$
$\Gamma_{coll,a,b}$ $=\gamma(\cos\theta+2\cos(2\theta)+2\cos(3\theta)+2\cos(4\theta)$
$\quad+\cos(5\theta)).$
$\delta\omega_{a}$ $=\gamma\sin(3\theta)+\gamma\sin(4\theta)+\frac{\gamma}{2}\sin\theta+\frac{\gamma}{2}\sin(7\theta),$ (16)
$\delta\omega_{b}$ $=\gamma\sin\theta+\gamma\sin(4\theta)+\frac{\gamma}{2}\sin(3\theta)+\frac{\gamma}{2}\sin(5\theta),$
$g_{a,b}$ $=\frac{\gamma}{2}(2\sin\theta+3\sin(2\theta)+\sin(3\theta)+\sin(5\theta)$
$\quad+\sin(6\theta)),$
$\Gamma_{a}$ $=2\gamma+2\gamma\cos(3\theta)+2\gamma\cos(4\theta)+\gamma\cos\theta$
$\quad+\gamma\cos(7\theta),$
$\Gamma_{b}$ $=2\gamma+2\gamma\cos\theta+2\gamma\cos(4\theta)+\gamma\cos(3\theta)$
$\quad+\gamma\cos(5\theta),$
$\Gamma_{coll,a,b}$ $=\gamma(2\cos\theta+3\cos(2\theta)+\cos(3\theta)+\cos(5\theta)$
$\quad+\cos(6\theta)).$
Using a strategy similar to that used in the braided configuration, we conclude that the dependence of exchange interaction and decay rates on $\theta$ in Eqs. (15) and (16) is a $2\pi$-periodic function. Therefore, we restrict our discussion to the region $\theta\in[0,2\pi]$, which fully captures the behavior of the system. In Fig. 4(b)(b) and Fig. 5(b)(b) we show the exchange interaction and decay rates from Eqs. (15) and (16) as a function of $\theta/\pi$, where $\theta\in[0,2\pi]$.
For the separate configuration, we observe from Fig. 4(b)(b) that there is no value of $\theta$ that satisfies the DFI condition (5). However, for the nested configuration, we note that the DFI condition (5) is satisfied when $\theta=\pi/4,3\pi/4,5\pi/4,7\pi/4$, as shown in Fig. 5(b)(b). The result demonstrates that our scheme enables DFI in the nested configuration, challenging the prevailing notion that DFI can only be achieved in the braided configuration [25, 26, 27, 28, 29].
@@
The concurrences $C_{s}$ and $C_{N}$ for two giant atoms in separate and nested configurations as a function of $\theta/\pi$ and $\gamma t$ are plotted in Fig. 4(c)(c) and Fig. 5(c)(c), respectively. From Fig. 4(c)(c) and Fig. 5(c)(c), we observe that both concurrences $C_{s}$ and $C_{N}$ reach a steady-state value at $\theta=0$ and $2\pi$, but drop to zero at $\theta=\pi$. These behaviors are consistent with those observed in the infinite waveguide scheme [30].

### v2

# Decoherence-free interaction and maximally entangled state generation in giant-atom semi-infinite waveguide systems

@@ Abstract
Giant atoms are artificial atoms that can couple to a waveguide non-locally. Previous works have shown that two giant atoms in a braided configuration can interact through one-dimensional (1D) infinite and chiral waveguides, with both individual and collective atomic relaxation being fully suppressed. In this paper, however, we show that the decoherence-free interaction (DFI) between two giant atoms can be realized in both braided and nested configurations when the waveguide is semi-infinite. This protected interaction fails to appear in semi-infinite waveguide systems containing two separate giant atoms or two small atoms. We also study the entanglement generation between two giant atoms coupled to a 1D semi-infinite waveguide. The results show that the maximally entangled state is generated in both braided and nested configurations due to the formation of DFI, and in the separate configuration, the maximally achievable entanglement can exceed 0.5. Finally, we generalize the discussion on DFI and entanglement generation to the case involving multiple giant atoms coupled into a semi-infinite waveguide. This study presents a new scheme for realizing DFI and generating maximally entangled states in giant-atom waveguide-QED systems.
@@ I INTRODUCTION
Recently, waveguide-QED systems containing giant atoms have garnered significant attention from many peers. In such systems, giant atoms are artificial atoms that can interact with the waveguide at multiple connection points [17]. Several interesting quantum optical phenomena have been demonstrated in giant-atom waveguide-QED systems, including frequency-dependent relaxation rates and Lamb shifts [18], non-Markovian decay dynamics [19, 20, 21], single-photon scattering [22, 23, 24], and DFI [25, 26, 27, 28, 29, 30, 31, 32]. In addition, the generation [33, 34, 35, 36, 37] and transfer [38] of quantum entanglement have also been investigated in giant-atom waveguide-QED systems.
However, it is important to note that in most studies on giant atoms, the waveguide is usually considered infinite or chiral. A key question is how giant atoms behave when coupled to a semi-infinite waveguide. The semi-infinite waveguide is a waveguide with one end terminated by a perfect mirror [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. Compared to the infinite and chiral waveguides, the semi-infinite waveguide is particularly notable in that the mirror at the end allows the atom to interact twice with the field in the waveguide. Hence, when the giant atom couples to the semi-infinite waveguide, its behavior may differ from that in the infinite and chiral waveguides. While the semi-infinite waveguide setup containing giant atoms has been addressed in previous studies [18, 52, 53], only a single giant atom was considered.
In this work, our study goes beyond the framework of a single giant atom coupled to a semi-infinite waveguide. We consider an arbitrary number of giant atoms $M$ interacting with a 1D semi-infinite waveguide. Treating the semi-infinite waveguide as the environment of atoms, we derive the Markov quantum master equation governing the dynamical evolution of $M$ giant atoms.
We focus on the case of two giant atoms. Concretely, we analyze the behavior of two giant atoms coupled to a 1D semi-infinite waveguide with three configurations: braided, separate, and nested. In doing so, we show that the DFI between two giant atoms, previously only observed within the braided configuration in 1D infinite [25, 29, 30, 31] and chiral [26, 27, 28] waveguide systems, can be realized in both braided and nested configurations in 1D semi-infinite waveguide systems. This protected interaction is not supported in a semi-infinite waveguide setup involving either two separate giant atoms or two small atoms. For two small atoms coupled to a semi-infinite waveguide, it has been reported that there is no DFI between them [25].
@@
Figure 1: (Color online) Two giant atoms in a braided configuration coupled to a 1D semi-infinite waveguide. The connection points are located at coordinates $x_{j1}$ and $x_{j2}$, where $j=a,b$. In our work, we consider the case where the two giant atoms share the same transition frequency, i.e., $\omega_{a}=\omega_{b}=\omega_{0}$.
@@
Finally, we generalize the discussion on DFI and entanglement generation to the case where multiple giant atoms are coupled to a semi-infinite waveguide. The results demonstrate that, for multiple giant atoms in both braided and nested configurations, DFI can still be realized, and a high degree of entanglement between the atoms can be generated.
@@
This paper begins with Sec. II, which describes the model for giant atoms coupled to a 1D semi-infinite waveguide. Then, we specifically explore the DFI and entanglement generation between two braided giant atoms in Sec. III and IV. In Sec. V, we consider the case of two giant atoms in the separate and nested configurations. In Sec. VI, we discuss how the results found here generalize to multiple atoms. Finally, we discuss the entanglement generation between two giant atoms in the two-excitation subspace and the experimental feasibility of the proposed scheme in Sec. VII, and summarize the main findings of this study in Sec. VIII. In addition, we provide four appendices. In Appendix A, the detailed derivation of the master equation is shown. In Appendix B, we present the analytical expressions for the concurrence that depend on the phase shift. In Appendix C, we replot the corresponding results for the infinite waveguide case, and relevant coefficients associated with separate and nested configurations in the semi-infinite waveguide system are provided in Appendix D.
@@ II THEORETICAL FRAMEWORK
We start by formulating a general model in which $M$ two-level giant atoms are coupled to a 1D semi-infinite waveguide along the positive $x$ semi-axis, with each atom $j$ having $N_{j}$ connection points. In this model, we assume the semi-infinite waveguide is realized by terminating the left end (coordinate set to $x=0$) of the 1D infinite waveguide with a mirror, and all atoms are located on the right side of the mirror.
@@
From Eq. (2), the primary difference with respect to the infinite [25, 29] and chiral [28] waveguide cases is that the expressions for $\delta\omega_{j}$, $g_{j,k}$, $\Gamma_{j}$, and $\Gamma_{coll,j,k}$ no longer depend just on the relative distances between the connection points: they are also determined by the distance from the connection points to the mirror in the semi-infinite waveguide case.
We focus on the case of two giant atoms where each atom interacts with the waveguide at two connection points. For two giant atoms with two connection points each, three distinct topology configurations are possible [25, 28]: braided, separate, and nested. The setup we initially consider is the braided configuration, as shown in Fig. 1. Later, we consider the separate and nested configurations in Sec. V. To simplify the discussion, the model in Fig. 1 adopts the following assumptions: $(i)$ the decay rate at each connection point is identical and denoted by $\gamma$, and $(ii)$ the phase shift between adjacent connection points is defined as $\theta$, where $\theta/2$ represents the phase shift accumulated as a photon travels from position $x_{a1}$ to the mirror ($x=0$). When the system contains two giant atoms, Eq. (1) reduces to
@@
$+\Gamma_{coll,a,b}\left[(\sigma_{-}^{a}\rho\sigma_{+}^{b}-\frac{1}{2}\left\{\sigma_{+}^{b}\sigma_{-}^{a},\rho\right\})+\text{H.c.}\right].$
@@
Figure 2: (Color online) Exchange interaction and decay rates as a function of $\theta/\pi$ for two braided giant atoms.
III DFI
In previous works, it has been confirmed that two braided giant atoms can interact via 1D infinite [25, 29, 30, 31] and chiral [26, 27, 28] waveguides without losing their excitations to the waveguide environment. That is, there is a DFI between two braided giant atoms. Mathematically, the condition for DFI to manifest can be stated as
@@ IV.1 Concurrence
In our work, we focus on entanglement generation in the case where only one giant atom is excited. When there is only one excitation in the system, we can remove the jump term (i.e., $\sigma_{-}^{a}\rho\sigma_{+}^{a},\sigma_{-}^{b}\rho\sigma_{+}^{b},\sigma_{-}^{a}\rho\sigma_{+}^{b}$, and $\sigma_{-}^{b}\rho\sigma_{+}^{a}$, which describe the effect of the measurement on the state of the system [72, 73, 74]) in the quantum master equation (3) to obtain a non-Hermitian effective Hamiltonian $H_{\text{eff}}$ [33, 37, 75, 76]:
${H}_{\text{eff}}$ $=\sum_{j=a,b}\delta\omega_{j}{\sigma}_{+}^{j}{\sigma}_{-}^{j}+g_{{a,b}}({\sigma}_{+}^{b}{\sigma}_{-}^{a}+{\sigma}_{+}^{a}{\sigma}_{-}^{b})$ (7)
$\quad-\frac{i}{2}\sum_{j=a,b}\Gamma_{j}{\sigma}_{+}^{j}{\sigma}_{-}^{j}-\frac{i}{2}\Gamma_{coll,a,b}({\sigma}_{+}^{b}{\sigma}_{-}^{a}+{\sigma}_{+}^{a}{\sigma}_{-}^{b}).$
@@
$i\frac{\partial}{\partial t}|\psi(t)\rangle={H}_{\text{eff}}|\psi(t)\rangle,$ (8)
@@ IV.2 Entanglement generation
$C_{eg}^{B}(t)=\frac{e^{-\frac{1}{2}\,\mathrm{Re}\left[R_{1}(t)\right]}}{2}\left|\frac{R_{2}(t)\,\left(R_{3}(t)+R_{4}(t)\right)}{AA^{*}}\right|.$ (12)
where the superscript $B$ and subscript $eg$ denotes the braided configuration case and atomic initial state $|e\rangle_{a}|g\rangle_{b}$, respectively. Similarly, superscripts $S$ and $N$, used in Sec. V, correspond to the separate and nested configurations. The expressions for $A$ and $R_{i}(t)$ ($i=1,2,3,4$), which depend on the phase shift $\theta$, are given in Appendix B.
@@
To clearly elucidate the phase-shift dependence of entanglement generation, we now analyze the evolution of the concurrence $C_{eg}^{B}$ when phase shift $\theta$ takes some specific values, where $\theta\in[0,\pi]$. As shown in Fig. 3(a)(a), when $\theta=0$, the concurrence $C_{eg}^{B}$ exhibits steady-state behaviors.
Figure 3: (Color online) $(a)$ Concurrence $C_{eg}^{B}$ as a function of $\theta/\pi$ and $\gamma t$. $(b)$ Concurrence $C_{eg}^{B}(t)$ as a function of $\gamma t$ at a given value of phase shift.
The steady-state behavior of the concurrence $C_{eg}^{B}$ at $\theta=0$ can be explained through the analytical expression. When $\theta=0$, the concurrence in Eq. (12) becomes
$C_{eg}^{B}(t)=\frac{1}{2}\left({1-e^{-16\gamma t}}\right).$ (13)
@@
$C_{eg}^{B}(t)=|\sin(2\gamma t)|.$ (14)
According to Eq. (14), it can be found that the concurrence $C_{eg}^{B}(t)$ oscillates between zero and one with a period $\frac{\pi}{2\gamma}$. When $t=\frac{n\pi}{4\gamma}$ (where $n$ is an odd integer), concurrence $C_{eg}^{B}(t)=1$, indicating that the two giant atoms reach a maximally entangled state. Specifically, when $\theta$ is taken to be $\pi$, no entanglement is generated here, i.e., $C_{eg}^{B}=0$. The reason is mainly that when $\theta=\pi$, exchange interaction $g_{a,b}$, individual decay rates $\Gamma_{a}$ and $\Gamma_{b}$, as well as the collective decay rate $\Gamma_{\text{coll},a,b}$ all become zero (as can be seen from Fig. 2), which implies that the two giant atoms decouple from the waveguide, thereby making entanglement generation impossible.
Figure 4: (Color online) $(a)$ Two giant atoms in a separate configuration coupled to a 1D semi-infinite waveguide. $(b)$ Exchange interaction and decay rates as a function of $\theta/\pi$ for two separate giant atoms. $(c)$ Concurrence $C_{eg}^{S}$ as a function of $\theta/\pi$ and $\gamma t$.
Figure 5: (Color online) $(a)$ Two giant atoms in a nested configuration coupled to a 1D semi-infinite waveguide. $(b)$ Exchange interaction and decay rates as a function of $\theta/\pi$ for two nested giant atoms. $(c)$ Concurrence $C_{eg}^{N}$ as a function of $\theta/\pi$ and $\gamma t$.
@@ V TWO GIANT ATOMS IN SEPARATE AND NESTED CONFIGURATIONS
In the above sections, we have analyzed the behavior of two braided giant atoms coupled to a semi-infinite waveguide, as well as their entanglement properties. In this section, as shown in Fig. 4 and Fig. 5, we consider the separate and nested configurations. According to Eq. (2), the relevant coefficients in master equation (3) for separate [Fig. 4(a)(a)] and nested [Fig. 5(a)(a)] configurations can be obtained. For the sake of conciseness, these coefficients are provided in Table 1 in Appendix D.
Following the strategy analogous to that used in the braided configuration, we conclude that the exchange interaction and decay rates for two giant atoms in separate and nested configurations are $2\pi$-periodic functions of $\theta$. Therefore, we restrict our discussion to the region $\theta\in[0,2\pi]$, which fully captures the behavior of the system. In Fig. 4(b)(b) and Fig. 5(b)(b) we show the exchange interaction and decay rates from Table 1 in Appendix D as a function of $\theta/\pi$, where $\theta\in[0,2\pi]$.
For the separate configuration, we observe from Fig. 4(b)(b) that there is no value of $\theta$ that satisfies the DFI condition (5). However, for the nested configuration, we note that the DFI condition (5) is satisfied when $\theta=\pi/4,3\pi/4,5\pi/4,7\pi/4$, as shown in Fig. 5(b)(b). The result demonstrates that our scheme enables DFI in the nested configuration, challenging the prevailing notion that DFI can only be achieved in the braided configuration [25, 26, 27, 28, 29, 30, 31, 32].
@@
The concurrences $C_{eg}^{S}$ and $C_{eg}^{N}$ for two giant atoms in separate and nested configurations as a function of $\theta/\pi$ and $\gamma t$ are plotted in Fig. 4(c)(c) and Fig. 5(c)(c), respectively. From Fig. 4(c)(c) and Fig. 5(c)(c), we observe that both concurrences $C_{eg}^{S}$ and $C_{eg}^{N}$ reach a steady-state value at $\theta=0$ and $2\pi$, but drop to zero at $\theta=\pi$. These behaviors are consistent with those observed in the infinite waveguide scheme [33], as presented in Appendix C.